15. How many moles of air are there in a 2.0 L bottle at 19ºC and 747 mm Hg?


1
Expert's answer
2021-04-23T08:12:13-0400

PV=nRT


R = 8.314 J/K mol

T=19+273=292k

V=2.0L

P=747mm Hg


n= moles =

"Moles = \frac{747×2}{8.314×292}=
0.615 moles"